A car increases its speed from 30 mph to 60 mph in 10 seconds. (a) Determine your acceleration,
in m / s2
, assuming it is constant. (b) What distance, in meters, does it travel in that time?
Answer #1

Solution) u = 30 mph

1 mph = 0.447 m/s

u = 30×0.447 = 13.41 m/s

v = 60 mph = 60×0.447 = 26.82 m/s

t = 10 s

(a) Acceleration , a = ?

v = u + at

a = ( v - u )/(t)

a = (26.82 - 13.41)/(10)

a = 1.341 m/s^2

(b) Distance, S = ?

v^2 = u^2 + 2aS

S = ( v^2 - u^2 )/(2a)

S = (26.82^2 - 13.41^2)/(2×1.341)

S = 201.15 m